OpenStudy (anonymous):

what is the solution of: d^2y/dx^2 + dy/dx - 6y=0

OpenStudy (anonymous):

\[\frac{ d ^{2}y }{dx ^{2} }+ \frac{ dy }{ dx }-6y =0\]

OpenStudy (anonymous):

well...have u tried it yet? setting up charecterestic equation and..

OpenStudy (anonymous):

i tried it but i couldnt get the solution

OpenStudy (anonymous):

what is the charecterestic equation here?

OpenStudy (anonymous):

@mukushla

OpenStudy (anonymous):

umm.. i'd probably say that's incorrect

OpenStudy (anonymous):

u just need to set up charecterestic equation\[m^2+m-6=0\]\[(m+3)(m-2)=0\]and then the answer will be\[y=c_1e^{m_1x}+c_2e^{m_2x}\]are u familiar with this type of solving ODE's?

OpenStudy (anonymous):

what is wrong in my solution